Circular No. 7838

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2000ft IN NGC 7469 L. Colina, Instituto de Física de Cantabria, Consejo Superior de Inves-

tigaciones Científicas (CSIC); A. Alberdi, Instituto de Astrofísica de Andalucía, CSIC; J. M. Torrelles, Institut d’Estudis Espacials de Catalunya, CSIC; N. Panagia, European Space Agency and Space Telescope Science Institute; A. S. Wilson, University of Maryland; and S. T. Garrington, Jodrell Bank Observatory, communicate: “The source RSN J230315+0852.4 in NGC 7469, first detected at radio frequencies on 2000 Oct. 27 (see IAUC 7587) and identified with a radio supernova (Colina et al. 2001, Ap.J. 553, L19), has been monitored extensively over the past few months with MER-LIN. The flux measured at 5 GHz for the last epoch (2001 Sept. 11) was 0.79 ± 0.02 mJy, indicating that the radio flux emitted by this source has dropped by about a factor of 2 since first discovered. Additional new data taken at 8.4 GHz with the Very Large Array (VLA) on 2002 Feb. 8 indicate that the decline in flux continues. The observed variation in flux is typical of other radio supernovae also identified as such in the optical, and confirms the source RSN J230315+0852.4 as a radio supernova (and the first not detected in the optical). Radio monitoring with MERLIN and the VLA is continuing.”

SUPERNOVA 2002ap IN M74 J. Danziger, Trieste Observatory; M. Della Valle, Arcetri Observatory;

E. Palazzi, Consiglio Nazionale delle Ricerche, Bologna; and E. Pian, Trieste Observatory, report that high-quality spectropolarimetry in the infrared bands J, H, and K were obtained with the 3.6-m reflector (+ NICS) by the Telescopio Nazionale Galileo staff at La Palma during inferior weather conditions on Feb. 6.9 UT. Polarization measures in the range 1.0– 2.4 μm fall in the range 0.5–0.9 percent, but the 3σ error bars suggest this could range from ∼ 0 to ∼ 2 percent. The spectra reveal a P-Cyg absorption feature near 1.0257 μm, which, if attributed solely to He I 1.083-μm, leads to a velocity of 16 500 km/s. Velocities of the blue-shifted absorption extend to 23 500 km/s. Another broad absorption feature is measured at 1.5136 μm. There appears to be a real abrupt drop in the level of the spectrum between the H and K spectral regions. Comparison with spectra of SN 1998bw (Patat et al. 2001, Ap.J. 555, 900) reveals a similarity. This apparent jump might plausibly be attributed to the presence of (a) broad emission feature(s) spanning the range 1.5–1.9 μm.

2002 February 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green